OpenStudy (anonymous):

9n-10>-82 and -1-2n> (or equal to)-5 solve and graph the solution

OpenStudy (espex):

Okay, and your attempt at this problem?

OpenStudy (anonymous):

Let's just say I need a ton of help

OpenStudy (espex):

Since these only have one variable, are we solving each individually and then graphing each of them?

OpenStudy (anonymous):

Correct and I have no clue how to do so

OpenStudy (espex):

Take the first one, 9n-10>-82, treat this as you would any other algebraic problem. How would you isolate n?

OpenStudy (anonymous):

Add 82 to both sides

OpenStudy (espex):

Not the way I would start but okay, so you have 9n+72>0, now what? And remember you want 'n' all alone.

OpenStudy (anonymous):

Aw so it'll take longer if I add 82 so let's try adding 10 to both sides

OpenStudy (espex):

Okay, so you now have 9n>-72, what's next?

OpenStudy (anonymous):

I would divide by 9

OpenStudy (espex):

So n> than ?

OpenStudy (anonymous):

7

OpenStudy (espex):

In the last problem you proved that 9*7 was 63.

OpenStudy (anonymous):

I meant 8

OpenStudy (espex):

Okay, and you have a -72

OpenStudy (anonymous):

So what's next

OpenStudy (espex):

Well first you need to recognize that if 9*8=72, then 9*(-8)=-72 and so your answer is n>(-8)

OpenStudy (anonymous):

That's it

OpenStudy (espex):

What this tells you is that the statement, 9n-10>-82, is true for all values of n>(-8)

OpenStudy (espex):

So when you graph this you will start with a circle around -8, because it is values > and not >=, and that your line will extend to all values that are greater than -8.

OpenStudy (anonymous):

Alright thanks man you've been a great help

OpenStudy (anonymous):

One more question is it an open or a closed circle

OpenStudy (espex):

You're welcome. Though you now need to do the same process for your -1-2n>= -5

OpenStudy (espex):

What does the closed circle represent?

OpenStudy (anonymous):

The answer is included on the graph

OpenStudy (espex):

I guess I should have asked if you were doing shading as well.

OpenStudy (espex):

So to include something, we would have to had that as an answer in the inequality. We showed that n > -8, and so it is no included in the answer.

OpenStudy (espex):

Some textbooks have a solid dot indicate inclusion, some have a full circle indicate above and below but not the exact value, and some have a half circle show that it is only from one side.

OpenStudy (anonymous):

So for the -8 it's a closed circle correct

OpenStudy (espex):

I would draw it like this, because we are only interested in values greater:

OpenStudy (espex):

But look at the examples your book gives and adjust your answer to match the lesson they have.

OpenStudy (anonymous):

Alright will do

OpenStudy (espex):

Good luck.